OpenStudy (anonymous):

Calculate the pH for a 0.01 M HCl solution without using a calculator

OpenStudy (sweetburger):

\[pH=- \log [H3O^+]\]

OpenStudy (sweetburger):

remember the greater the concentration the lower the pH

OpenStudy (anonymous):

Knowing that HCl is a strong acid, and strong acids almost completely dissociate in water\[[HCl] = [H_3O^+]\] Therefore \[pH = -\log[H_3O^+] = -\log[HCl] = -\log(0.01) = -\log(10^{-3})\] This is where you need to know the basics of logarithms, log (as we've written it) is to the base ten, this means that the log of 10 to any power will equal that power. Remember to account for the negative sign too

OpenStudy (sweetburger):

10^-?=.01

OpenStudy (sweetburger):

u can figure the variable out

OpenStudy (anonymous):

ok i got it thx